Suppose a firm produces 15 units of output per month and has a total variable cost of R5 000 per month. If its average fixed costs are R350,what are the total coat per month?
1
Expert's answer
2020-10-01T10:01:51-0400

"Total\ Costs= Total Fixed Costs+Total\ variable\ Cost\\
But\ Average\ Variable\ Fixed\ Costs=\frac{Total Fixed\ cost}{Quantity}\\
\implies\ Total\ fixed\ costs=Quantity\times\ Average\ variable\ fixed\ costs\\
=15\times350=5,250\\
Total\ Costs
=5,250+5,000\\
=R10,250"